# Generalised Causal Dantzig

### v1

@@ 2.2 Generalised causal Dantzig characterization of causal model
Uniqueness is achieved under the restriction that $f_{S}$ depends on all $\boldsymbol{x}_{S}$ — in Theorem 3 this was referred to as strict dependence on $\boldsymbol{x}$. Any set $\boldsymbol{X}_{S}$ that contains $\boldsymbol{X}_{PA}$ together with variables that are d-separated from $Y$ by its parents would still satisfy the two properties that characterise the causal model. This is because $f_{S}$ is simply constant with respect to these variables.
@@
    For any set of covariates $\boldsymbol{X}_{S}$, with $S\subset\{1,\ldots,p\}$, determine the population maximum likelihood,
@@
    If it is, then $\boldsymbol{\beta}_{S}^{\top}\psi_{S}$ is a candidate for $f_{PA}=\boldsymbol{\beta}_{PA}^{\top}\psi_{PA}$.
@@
    Among the set of candidate models $\boldsymbol{\beta}_{S}$, select $\boldsymbol{\beta}_{PA}$ as the one minimising the Bayesian Information Criterion (BIC).
@@ 2.3 Population and empirical algorithms
Theorem 3 can be translated into a population algorithm for finding the causal model. It involves a combinatorial search among all subsets of covariates, and is described in Algorithm 1. The algorithm involves a search of the models that achieve a perfectly dispersed Pearson risk. Then, an information criterion is used to select the optimal model among the class of candidate models. The choice of an information criterion is motivated by the fact that variables that are d-separated from $Y$ by its causal parents, such as ancestors of $Y$, have no predictive power on $Y$. In principle, any information criterion that penalises the expected likelihood by a term including the number of covariates can be used at this stage.
For a finite sample, $(\boldsymbol{x}_{i},y_{i})$, $i=1,\ldots,n$, an empirical version of the population algorithm is proposed in Algorithm 2. It involves replacing the population likelihood by the empirical likelihood, while the requirement of a perfectly dispersed Pearson risk is checked via an appropriate statistical test. In particular, for each candidate subset $\boldsymbol{X}_{S}$, we test the null hypothesis
$R^{+}=\displaystyle\sum_{i=1}^{n}\left[\frac{(y_{i}-\dot{b}({\widehat{\boldsymbol{\beta}}}_{S}^{\top}\tilde{\psi}_{S}(\boldsymbol{x}_{i,S})))^{2}}{\ddot{b}({{\widehat{\boldsymbol{\beta}}}_{S}}^{\top}\tilde{\psi}_{S}(\boldsymbol{x}_{i,S}))a(\phi)}\right],$
where $\boldsymbol{x}_{i,S}$ denotes the $i$-th realization of $\boldsymbol{X}_{S}$ and $\widehat{\boldsymbol{\beta}}_{S}$ is a penalised maximum likelihood estimator,
for some convenient smoothness or sparsity penalty $P_{\lambda}$ (Wood, 2017). Under the null hypothesis and some standard conditions on the penalty, the Pearson statistic is asymptotically chi-squared distributed. In particular, it holds that, approximately,
$R^{+}\overset{H_{0}}{\sim}\chi^{2}_{n-\hat{q}},$ (1)
with $\hat{q}$ an estimate for the effective degrees of freedom of the model.
@@
    For any set of covariates $\boldsymbol{X}_{S}$, with $S\subset\{1,\ldots,p\}$, find the penalised maximum likelihood estimator,
@@
    Find $S_{1},\ldots,S_{K}\subset\{1,\ldots,p\}$ such that
@@
    cannot be rejected, i.e., such that
@@
    for some significance level $\alpha$, where $\hat{q}_{k}$ is the effective degrees of freedom of the model involving $S_{k}$.
@@
    Among the set of candidate models $f_{S_{1}},\ldots,f_{S_{K}}$, select $f_{PA}$ as the one minimising the Bayesian Information Criterion (BIC).
Algorithm 2 Empirical algorithm for finding causal parents.
The algorithm considers all possible models and retains those for which the null hypothesis cannot be rejected. As the aim is to select perfectly dispersed models, a two-sided test is needed and therefore the acceptance region for the model that includes variables $S_{k}$ is given by
for some significance level $\alpha$, where $\hat{q}_{k}$ is the effective degrees of freedom associated with the model involving $S_{k}$. As with the population algorithm, the candidate set of models obtained via statistical testing is finally refined via BIC, in order to find the sparse most predictive model among the class of perfectly dispersed models.
Looking at all possible models leads to a computational complexity of the algorithm that is exponential. For real applications involving a large number of covariates, we therefore propose also a stepwise version of the algorithm. At each step, we add the variable with the largest p-value from the statistical test (1) and stop when the addition of the new variable decreases the p-value of the resulting model and drops below the significance level $\alpha$.
@@ 3 Simulation study
In this section, we illustrate the performance of the generalised causal Dantzig procedure. In section 3.1, we describe the data generating process for a generalised additive target inside a structural equation model. In section 3.2 we explain the idea behind Theorem 2 as the characterization of the causal parameters in the population scenario by means of a large sample. Finally in section 3.3 we show the performance of the method in a non-linear Poisson regression setting for the target of interest.
3.1 Simulating data via an underlying SEM
We generate a sample $(\boldsymbol{x}_{i},y_{i})$, $i=1,\ldots,n$ from a structural equation model with exponential target in definition 1 by first simulating $(\boldsymbol{x}_{i},z_{i})$ observations from a linear structural equation model (SEM), as follows,
$\begin{bmatrix}Z\\ \boldsymbol{X}\\ \end{bmatrix}=B\begin{bmatrix}Z\\ \boldsymbol{X}\\ \end{bmatrix}+\begin{bmatrix}\epsilon_{Z}\\ \boldsymbol{\epsilon}_{X}\\ \end{bmatrix}$ (2)
where $B=\begin{bmatrix}0&\boldsymbol{b}_{PA}^{\top}\\ \boldsymbol{\beta}_{CH}&B_{X}\\ \end{bmatrix}$ is the matrix of unknown parameters describing the structural equations between the variables $(\boldsymbol{X},Z)$, such that $(I-B)$ is invertible, and $\epsilon_{Z}\sim N(0,\sigma^{2}_{Z})$ and $\epsilon_{j,X}\sim N(0,\sigma_{j}^{2}))$ are independent noise components ($j=1,\ldots,p$). The exponential family target variables $Y_{i}$ are then simulated via
$y_{i}=F^{-1}_{b(\boldsymbol{\beta}_{PA}^{\top}\boldsymbol{x}_{i,PA}),a(\phi)}(\Phi_{0,\sigma^{2}_{Z}}(\epsilon_{i,Z})),$ (3)
where $F_{b(\boldsymbol{\beta}_{PA}^{\top}\boldsymbol{x}_{i,PA}),a(\phi)}$ is the cdf of the EDF distribution and $\Phi_{m,s^{2}}$ is the cdf of a $N(m,s^{2})$ distribution. In this notation, $\boldsymbol{\beta}_{PA}$ is the vector of non-zero coefficients corresponding to a causal generalised linear model for the target variable $Y$. The vector $\boldsymbol{b}_{PA}\in\mathbb{R}^{p}$ is identical to $\boldsymbol{\beta}_{PA}$, but also includes additional zeroes for the variables that are not the causal parents.
This formulation of the model allows to control the predictive power of the children of $Y$ via the underlying linear SEM (2), where the variable $Z$ is in the same scale as the predictors $\boldsymbol{X}$. In this way, we can make the detection of the causal parameters $\boldsymbol{\beta}_{PA}$ sufficiently different from the standard regression approaches. Figures 1a and 1c show two causal DAGs that will be used in the simulations in the next two sections. The left plots show the causal graphs of $(\boldsymbol{X},Y)$, while the right plots show the DAGs corresponding to the equivalent formulation as a linear SEM on $(\boldsymbol{X},Z)$. The transformation (3) guarantees that $Y$ has the required distribution, i.e.,
$Y|\boldsymbol{X}_{PA}=\boldsymbol{x}_{PA}\sim EDF\Big{(}b(\boldsymbol{\beta}_{PA}^{\top}\boldsymbol{x}_{PA}),a(\phi)\Big{)},$
with the parents of $Y$ being the parents of $Z$ in the SEM. In the simulation studies, we will consider Poisson regression as a running example.
@@ 3.1 Simulating data via an underlying SEM
[…]
Figure 1: Left: Causal graphical models for generating data on a target variable $Y$ and (a) p=2 (c) p=7 predictors. (b), (d): Equivalent formulations of (a),(c), respectively, with a latent variable $Z$, a linear SEM on $(\boldsymbol{X},Z)$ and a transformation of $Z$ in a $Y$ with a given EDF distribution conditional on $\boldsymbol{X}_{PA}$.
3.2 Proof-of-concept: population scenario
In this section, we highlight the main features of the proposed method via a simple simulation. We consider the DAG in Figure 1(a), with a target variable $Y$ having a causal parent $X_{1}$ and a child $X_{2}$. We consider a Poisson distribution for $Y$ conditional on $X_{1}$ and an underlying structural equation model (Figure 1(b)) given by
$\begin{cases}X_{1}&=\epsilon_{1}\\ Z&=X_{1}+\epsilon_{Z}\\ X_{2}&=Z+\epsilon_{2},\end{cases}$
where we assume the noise components $\epsilon_{1}$, $\epsilon_{2}$ and $\epsilon_{Z}$ to be independent and identically distributed standard normal random variables. From the structural equations, we deduce that $\boldsymbol{b}_{PA}=(1,0)^{\top}$, with the causal parameter $\beta_{PA}=1$, so
$Y|X_{PA}=x_{PA}\sim\mbox{Poisson}(e^{X_{1}}).$
@@ 3.2 Proof-of-concept: population scenario
For easy of visualization, we consider a full model without the intercept, and so a $\boldsymbol{\beta}$ vector of length 2. As expected, Figure 2(a) shows how the Maximum Likelihood (ML) estimate of $\boldsymbol{\beta}$ achieves the maximum of this function, being the most predictive model on the observed data. On the other hand, the causal parameter $\boldsymbol{b}_{PA}$ lies on a lower contour line.
Figure 2: Simulated data ($n=100000$) using a Poisson regression model with $p=2$ predictors. Contour lines of likelihood on (a) observational data and (b)-(c) two future environments, with covariates $X_{1}$ and $X_{2}$ perturbed by an additive noise with variance $\sigma^{2}=5$ and $\sigma^{2}=10$, respectively. The MLE solution (red dot) has a higher likelihood than the causal solution (blue dot) on the observational environment (a) but a lower likelihood on the shifted environments (b-c), particularly for larger noise (c). That is, the causal model has better out-of-sample guarantees. (d) Contour line for Pearson risk $=1$ contains the true $\boldsymbol{b}_{PA}$, which coincide with the ML $\boldsymbol{\beta}_{1}$ of the model that includes only the causal parent $X_{1}$. The ML population estimates $\boldsymbol{\beta}_{S}$ from other models are not on the Pearson risk $=1$ surface.
@@
Figure 2(d) shows the surface satisfied by all $\boldsymbol{\beta}$ with Pearson risk equal to 1. As expected from Theorem 2, $\boldsymbol{b}_{PA}$ belongs to this surface, and so does the $\boldsymbol{\beta}_{1}$, the population maximum likelihood estimator of the dependence of $Y$ only on the causal parent $X_{1}$. The remaining two models, i.e., the saturated model ($\beta_{12}$) and the model with $Y$ dependent only on $X_{2}$ ($\beta_{2}$), lead to population maximum likelihood estimates that do not lie on the Pearson risk surface. Indeed, these models are not causal and, therefore, do not enjoy the invariance property, as shown in the uniqueness result of Theorem 3.
3.3 Finite sample behaviour of generalised causal Dantzig
@@
In this section, we study the behaviour of the finite sample algorithm to obtain the causal parents of target $Y$. We consider the more complex system in Figure 1(c), where $Y$ has 2 causal parents ($X_{2}$ and $X_{3}$) and with 5 other variables in the system. We simulate it according to a latent structural equation model shown in Figure 1(d), where
$\begin{cases}X_{1}&=\epsilon_{1}\\ X_{2}&=X_{1}+\epsilon_{2}\\ X_{3}&=X_{1}+X_{2}+\epsilon_{3}\\ Z&=\sin(5X_{2})+X_{3}^{3}+\epsilon_{Z}\\ X_{4}&=X_{2}+\epsilon_{4}\\ X_{5}&=Z+\epsilon_{5}\\ X_{6}&=Z+\epsilon_{6}\\ X_{7}&=X_{6}+\epsilon_{7}\\ \end{cases}.$
@@
In particular, the non-linear predictor for the target $Y$ is described by the function $f_{PA}(x_{2},x_{3})=\sin(5x_{2})+x_{3}^{3}$. The error terms are assumed to be independent and normally distributed with mean 0 and variance 0.04. For the distribution of $Y$ given $X_{2}$ and $X_{3}$, we consider a Poisson regression model, i.e.,
$Y|\boldsymbol{X}_{PA}=\boldsymbol{x}_{PA}\sim\mbox{Poisson}\Big{(}e^{f_{PA}(\boldsymbol{x}_{PA})}\Big{)}.$
PC-algorithm
$n$ GCD $X_{2}\rightarrow Y\leftarrow X_{3}$ $X_{2}\rightarrow Y$ $Y\leftarrow X_{3}$
1000 91% 9% 24% 16%
500 80% 0% 14% 1%
250 60% 0% 11% 2%
200 49% 0% 9% 0%
150 27% 0% 2% 0%
100 8% 0% 1% 0%
@@
Figure 3: Results across 100 simulations with $n=1000$ from the Model in Figure 1(c) with a Poisson distribution for $Y$ conditional on its causal parents. (a) BIC values across all models show how the causal model (red vertical line) is not the one that minimises BIC; (b) Pearson risk values show how the causal model has a median Pearson risk close to 1; (c) % times a model is selected by a Generalised Causal Dantzig approach using a 5% significance level shows a high percentage for the causal model; (d)The table reports the percentage of times that the correct model is selected by the proposed Generalised Causal Dantzig (GCD) approach. For the PC-algorithm, we count the graphs where there is a direct arrow to the target $Y$ from its parents ($X_{2}$, $X_{3}$), as well as those where only one of the two arrows is detected. Across all settings, the proposed method detects the true causal model more often than the PC-algorithm.
@@
We repeat the simulation 100 times with sample size $n=1000$. We fit Poisson regression models for $Y$ for all possible subsets of the predictors. In order not to use any information from the generative process, we include also an intercept term in each of these models. Figure 3(a) shows the boxplots of BIC values of the models across the 100 simulations, sorted in descending order according to their median. The figure shows how the causal model (red vertical line) is far from being the most predictive model on the observational data. Indeed, the model with the lowest median BIC value is a model that includes the parents of $Y$ ($X_{2}$ and $X_{3}$) as well as its children ($X_{5}$ and $X_{6}$), which form the Markov blanket of $Y$ in the graph in Figure 1(c). On the other hand, Figure 3(b) shows how the causal model achieves a median Pearson risk close to 1, together with 3 other models that contain either ancestors ($X_{1}$) or variables blocked by the parents of $Y$ ($X_{4}$). Figure 3(c) shows how the generalised causal Dantzig procedure is able to detect the correct causal model 91% of the time out of all models, when setting an $\alpha=5\%$ in the statistical testing procedure.
@@
Figure 3(d) shows how the detecting accuracy changes for decreasing sample sizes $n$. As expected, the accuracy in detecting the causal model decreases with $n$. However, it is significantly larger than what can be achieved with the PC-algorithm. As there are no implementations of the PC-algorithm that are specific to systems with Gaussian variables and a target count variable, we stabilised the variance of $Y$ by taking a square-root transformation and constructed the causal graph using Gaussian conditional independence tests on the resulting data.
4 Identifying causal drivers in two empirical studies

### v2

@@ 2.2 Characterization of causal model
Uniqueness is achieved under the restriction that $f_{S}$ depends on all $\boldsymbol{x}_{S}$ — in Theorem 3 this was referred to as strict dependence on $\boldsymbol{x}$. Any set $\boldsymbol{X}_{S}$ that contains $\boldsymbol{X}_{\mbox{\scriptsize PA}}$ together with variables that are d-separated from $Y$ by its parents would still satisfy the two properties that characterize the causal model. This is because $f_{S}$ is simply constant with respect to these variables.
@@
    For any set of covariates $\boldsymbol{X}_{S}$, with $S\subset\{1,\ldots,p\}$, determine the maximum of the expected likelihood,
@@
    If it is, then $\boldsymbol{\beta}_{S}^{\top}\psi_{S}$ is a candidate for $f_{\mbox{\scriptsize PA}}=\boldsymbol{\beta}_{\mbox{\scriptsize PA}}^{\top}\psi_{\mbox{\scriptsize PA}}$.
@@
    Among the set of candidate models $\boldsymbol{\beta}_{S}$, select $\boldsymbol{\beta}_{\mbox{\scriptsize PA}}$ as the one minimizing the Bayesian Information Criterion (BIC).
@@ 2.3 Population and empirical algorithms
Theorem 3 can be translated into a population algorithm for finding the causal model. This involves a combinatorial search among all subsets of covariates, and is described in Algorithm 1. The algorithm involves a search of the models that achieve a perfectly dispersed Pearson risk. Then, an information criterion is used to select the optimal model among the class of candidate models. The choice of an information criterion is motivated by the fact that variables that are d-separated from $Y$ by its causal parents, such as ancestors of $Y$, have no predictive power on $Y$. In principle, any information criterion that penalizes the expected likelihood by a term including the number of covariates can be used at this stage.
@@
    For any set of covariates $\boldsymbol{X}_{S}$, with $S\subset\{1,\ldots,p\}$, find the penalized maximum likelihood estimator,
@@
    Find $S_{1},\ldots,S_{K}\subset\{1,\ldots,p\}$ such that for $k=1,\ldots,K$
@@
    cannot be rejected for some significance level $\alpha$.
  3. 3.
    Among the set of candidate models $f_{S_{1}},\ldots,f_{S_{K}}$, select $f_{\mbox{\scriptsize PA}}$ as the one minimising the Bayesian Information Criterion (BIC).
Algorithm 2 Full empirical algorithm for finding causal parents.
For a finite sample, $(\boldsymbol{x}_{i},y_{i})$, $i=1,\ldots,n$, an empirical version of the population algorithm is proposed in Algorithm 2. This involves replacing the expected likelihood by the empirical likelihood, while the requirement of a perfectly dispersed Pearson risk is checked via an appropriate statistical test. In particular, for each candidate subset $\boldsymbol{X}_{S}$, we test the null hypothesis
$R^{+}_{S}=\displaystyle\sum_{i=1}^{n}\left[\frac{(y_{i}-\dot{b}({\widehat{\boldsymbol{\beta}}}_{S}^{\top}\tilde{\psi}_{S}(\boldsymbol{x}_{i,S})))^{2}}{\ddot{b}({{\widehat{\boldsymbol{\beta}}}_{S}}^{\top}\tilde{\psi}_{S}(\boldsymbol{x}_{i,S}))a(\phi)}\right],$
where $\boldsymbol{x}_{i,S}$ denotes the $i$-th realization of $\boldsymbol{X}_{S}$ and $\widehat{\boldsymbol{\beta}}_{S}$ is a penalized maximum likelihood estimator,
for some convenient smoothness or sparsity penalty $P_{\lambda}$ (Wood, 2017).
The distribution of the test statistic under the null hypothesis is generally unknown, so bootstrap can be used to assess the evidence against the null hypothesis. The algorithm considers all possible models and retains those for which the null hypothesis cannot be rejected. Since the aim is to select perfectly dispersed models, a two-sided test is needed. As with the population algorithm, the candidate set of models obtained via statistical testing is finally refined via BIC, in order to find the sparse most predictive model among the class of perfectly dispersed models.
For some generalized linear models, an approximate distribution of the test statistic under the null hypothesis is available, leading to computational efficiency. Of special notice is the case of Poisson regression: under the null hypothesis and some standard conditions on the penalty, the Pearson statistic can be shown to be asymptotically chi-squared distributed. In particular, it holds, approximately, for the sum of the squared Pearson residuals that
$R^{+}_{S}\overset{H_{0}}{\sim}\chi^{2}_{n-\hat{q}(S)},$ (1)
with $\hat{q}(S)$ an estimate for the effective degrees of freedom of the model associate with variables $S$. Therefore, in this case, the acceptance region for the model that includes variables $S_{k}$ can be efficiently approximated by
@@
for some significance level $\alpha$, where $\hat{q}_{k}$ is the effective degrees of freedom associated with the model involving $S_{k}$.
2.4 Stepwise algorithm
The full algorithm considers all possible, i.e., $2^{p}$, models, which leads to an exponential computational complexity of the algorithm. For applications involving a large number of covariates, this may be prohibitively slow. We therefore propose also a stepwise version of the algorithm. The algorithm is is described in detail in Algorithm 3 and is split into two phases. Starting from an empty model, in the first phase we add variables one-by-one until we get the largest model that we cannot reject to be perfectly dispersed. In a second phase, we remove superfluous, hopefully d-separated, variables by means of a BIC procedure.
  1. 1.
    Select a threshold $\alpha$.
  2. 2.
    Start with an empty model involving just the intercept;
@@ 2.4 Stepwise algorithm
    Calculate the p-value $p$ associated with the test of the perfect dispersion of the squared Pearson residuals.
  4. 4.
    Iterate
    1. (a)
      For all available non-included variables $X_{j}$, fit the model with that variable included to the current model.
    2. (b)
      Calculate p-value $p_{j}$ for the test of perfect dispersion associated with addition of variable $X_{j}$ to the existing model.
      • Add variable $X_{k}$ to the existing model, if $p_{k}=\max_{j}p_{j}$ and $p_{k}>p$ or $p_{k}>\alpha$. Continue with the iteration, setting $p=p_{k}$.
      • Stop with the iteration if $\max p_{j}<p$ and $\max p_{j}<\alpha$, and pass current model to step 5.
  5. 5.
    Starting from the model selected in step 4 with BIC $b$, iterate:
    1. (a)
      For all variables $X_{j}$ included in the current model, fit the model $M_{-j}$ without that variable $X_{j}$.
    2. (b)
      Calculate BIC $b_{j}$ for model $M_{-j}$ without variable $X_{j}$.
      • Remove variable $X_{k}$ from the existing model, if $b_{k}=\min_{j}b_{j}$ and $b_{k}<b$. Continue with the iteration, setting $b=b_{k}$.
      • Stop with the iteration, if $\min_{j}b_{j}>b$ and return the current model.
Algorithm 3 Stepwise algorithm for finding causal parents.
@@ 3 Simulation study
In this section, we illustrate the performance of the procedure. In section 3.1 we explain the idea behind Theorem 2 as the characterization of the causal parameters in the population scenario by means of a large sample. In the following two sections, we show the performance of the method in two important cases where causal discovery can be achieved from a single observational environment. In particular, in section 3.2 we consider a Poisson regression setting for the target of interest, while in section 3.3 we consider the case of logistic regression.
3.1 Proof-of-concept: population scenario
In this section, we highlight the main features of the proposed method via a small example. We consider the DAG in Figure 1, with a target variable $Y$ having a causal parent $X_{1}$ and a child $X_{2}$.
@@ 3.1 Proof-of-concept: population scenario
$\begin{cases}X_{1}&=\epsilon_{1}\\ Z&=X_{1}+\epsilon_{Z},\\ Y&=F_{\rm{PA}}^{-1}(\Phi_{0,\sigma^{2}_{Z}}(\epsilon_{Z}))\\ X_{2}&=Z+\epsilon_{2},\end{cases}$
Figure 1: Structural causal model for generating data on a target variable $Y$, $p=2$ predictors and the causal graph (a). The structural equations in (b) via a latent variable $Z$ show how the target $Y$ has a $\mbox{Poisson}(e^{X_{1}})$ distribution $F_{\rm{PA}}$ conditional on the causal parent $X_{1}$.
The formulation of the structural equations via a latent variable $Z$ (Figure 1(b)) allows to control more easily the predictive power of the children of $Y$, since this variable is in the same scale as the predictors $\boldsymbol{X}$. In this way, we can make the detection of the causal parameters $\boldsymbol{\beta}_{\mbox{\scriptsize PA}}$ sufficiently challenging. For the simulation, we assume the noise components $\epsilon_{1}$, $\epsilon_{2}$ and $\epsilon_{Z}$ to be independent and identically distributed standard normal random variables, while the target $Y$ has a Poisson distribution $F_{\rm{PA}}$ conditional on its causal parent $X_{1}$, given by
$Y|\boldsymbol{X}_{\mbox{\scriptsize PA}}=\boldsymbol{x}_{\mbox{\scriptsize PA}}\sim\mbox{Poisson}(e^{x_{1}}).$
@@
For easy of visualization, we consider a full model without the intercept, and so a $\boldsymbol{\beta}$ vector of length 2. As expected, Figure 2(a) shows how the Maximum Likelihood (ML) estimate of $\boldsymbol{\beta}$ achieves the maximum of this function, being the most predictive model on the observed data. On the other hand, the causal parameter $(1,0)$ lies on a lower contour line.
Figure 2: Simulated data ($n=100000$) using a Poisson regression model with $p=2$ predictors. Contour lines of likelihood on (a) observational data and (b)-(c) two future environments, with covariates $X_{1}$ and $X_{2}$ perturbed by an additive noise with variance $\sigma^{2}=5$ and $\sigma^{2}=10$, respectively. The MLE solution (red dot) has a higher likelihood than the causal solution (blue dot) on the observational environment (a) but a lower likelihood on the shifted environments (b-c), particularly for larger noise (c). That is, the causal model has better out-of-sample guarantees. (d) Contour line for Pearson risk $=1$ contains the true causal parameters, which coincide with the ML $\boldsymbol{\beta}_{1}$ of the model that includes only the causal parent $X_{1}$. The ML population estimates $\boldsymbol{\beta}_{S}$ from other models are not on the Pearson risk $=1$ surface.
@@
Figure 2(d) shows the surface satisfied by all $\boldsymbol{\beta}$ with Pearson risk equal to 1. As expected from Theorem 2, $\boldsymbol{b}_{\mbox{\scriptsize PA}}=(1,0)$ belongs to this surface, and so does $\boldsymbol{\beta}_{1}$, the population maximum likelihood estimator of the dependence of $Y$ only on the causal parent $X_{1}$. The remaining two models, i.e., the saturated model ($\beta_{12}$) and the model with $Y$ dependent only on $X_{2}$ ($\beta_{2}$), lead to population maximum likelihood estimates that do not lie on the Pearson risk surface. Indeed, these models are not causal and, therefore, do not enjoy the invariance property, as shown in the uniqueness result of Theorem 3.
3.2 Finite sample behaviour: causal Poisson regression
@@
In this section, we study the behaviour of the finite sample algorithm to obtain the causal parents of target $Y$. We consider the more complex system in Figure 3, where $Y$ has 2 causal parents ($X_{2}$ and $X_{3}$) and with 5 other variables in the system.
[…]
$\begin{cases}X_{1}&=\epsilon_{1}\\ X_{2}&=X_{1}+\epsilon_{2}\\ X_{3}&=X_{1}+X_{2}+\epsilon_{3}\\ Z&=\sin(5X_{2})+X_{3}^{3}+\epsilon_{Z}\\ Y&=F_{\rm{PA}}^{-1}(\Phi_{0,\sigma^{2}_{Z}}(\epsilon_{Z}))\\ X_{4}&=X_{2}+\epsilon_{4}\\ X_{5}&=Z+\epsilon_{5}\\ X_{6}&=Z+\epsilon_{6}\\ X_{7}&=X_{6}+\epsilon_{7},\end{cases}$
Figure 3: Structural causal model for generating data on a target variable $Y$, $p=7$ predictors and the causal graph (a). The structural equations in (b), via a latent variable $Z$, show how the target $Y$ has a $\mbox{Poisson}(\sin(5X_{2})+X_{3}^{3})$ distribution $F_{\rm{PA}}$ conditional on the causal parents $X_{2}$ and $X_{3}$.
The structural causal model in Figure 3(b) shows how the distribution $F_{\rm{PA}}$ of $Y$ given its causal parents $X_{2}$ and $X_{3}$, is
$Y|\boldsymbol{X}_{\mbox{\scriptsize PA}}=\boldsymbol{x}_{\mbox{\scriptsize PA}}\sim\mbox{Poisson}\Big{(}e^{\sin(5x_{2})+x_{3}^{3}}\Big{)},$
@@
that is, it is a Poisson regression where the predictor for the target $Y$ is described by the nonlinear function $f_{\mbox{\scriptsize PA}}(x_{2},x_{3})=\sin(5x_{2})+x_{3}^{3}$. The error terms are assumed to be independent and normally distributed with mean 0 and variance 0.04.
Causal GLM PC-algorithm
$n$ Full Step $X_{2},X_{3}$ $X_{2}$ $X_{3}$
1000 91% 60% 9% 24% 16%
500 80% 64% 0% 14% 1%
250 60% 49% 0% 11% 2%
200 49% 41% 0% 9% 0%
150 27% 33% 0% 2% 0%
100 8% 8% 0% 1% 0%
@@
Figure 4: Results across 100 simulations with $n=1000$ from the Model in Figure 3 with a Poisson distribution for $Y$ conditional on its causal parents. (a) BIC values across all models show how the causal model (red vertical line) is not the one that minimizes BIC; (b) Pearson risk values show how the causal model has a median Pearson risk close to 1; (c) % times a model is selected by a causal GLM approach using a 5% significance level shows a high percentage for the causal model; (d)The table reports the percentage of times that the correct model is selected by the proposed causal GLM approach, both under an exhaustive and a stepwise procedure. For the PC-algorithm, we count the graphs where there is a direct arrow to the target $Y$ from its parents ($X_{2}$, $X_{3}$), as well as those where only one of the two arrows is detected. Across all settings, the proposed method detects the true causal model more often than the PC-algorithm.
@@
We repeat the simulation 100 times with sample size $n=1000$. We fit Poisson regression models for $Y$ for all possible subsets of the predictors. In order not to use any information from the generative process, we include also an intercept term in each of these models. Figure 4(a) shows the boxplots of BIC values of the models across the 100 simulations, sorted in descending order according to their median. The figure shows how the causal model (red vertical line) is far from being the most predictive model on the observational data. Indeed, the model with the lowest median BIC value is a model that includes the parents of $Y$ ($X_{2}$ and $X_{3}$) as well as its children ($X_{5}$ and $X_{6}$), which form the Markov blanket of $Y$ in the graph in Figure 3. On the other hand, Figure 4(b) shows how the causal model achieves a median Pearson risk close to 1, together with 3 other models that contain either ancestors ($X_{1}$) or variables blocked by the parents of $Y$ ($X_{4}$). Figure 4(c) shows how the causal GLM procedure is able to detect the correct causal model 91% of the time out of all models, when setting an $\alpha=5\%$ in the asymptotic chi-squared statistical testing procedure.
@@
Figure 4(d) shows how the detection accuracy changes for decreasing sample sizes $n$. As expected, the accuracy in detecting the causal model decreases with $n$. Using a stepwise procedure rather than an exhaustive search generally leads to a reduction in the detection accuracy, but at a much lower computational cost (on average 5.2 times faster than the exhaustive search in this simulation study). In general, the accuracy achieved also by the stepwise procedure is significantly larger than what can be achieved with the PC-algorithm. As there are no implementations of the PC-algorithm that are specific to systems with Gaussian variables and a target count variable, we stabilized the variance of $Y$ by taking a square-root transformation and constructed the causal graph using Gaussian conditional independence tests on the resulting data.
3.3 Finite sample behaviour: causal logistic regression
In this section, we evaluate the performance of the method in the case of a binary target variable $Y$, whose relationship with its causal parents is described by a logistic regression model. We consider the system in Figure 5, where $Y$ has 2 causal parents ($X_{2}$ and $X_{3}$) and with 3 other variables in the system.
[…]
$\begin{cases}X_{1}&=\epsilon_{1}\\ X_{2}&=X_{1}+\epsilon_{2}\\ X_{3}&=\epsilon_{3}\\ Y&\sim\rm{Bernoulli}\Big{(}\dfrac{e^{0.8X_{2}-0.9X_{3}}}{1+e^{0.8X_{2}-0.9X_{3}}}\Big{)}\\ X_{4}&=X_{2}+\epsilon_{4}\\ X_{5}&=(1-\pi)Y+\pi(1-Y)+\epsilon_{5},\end{cases}$
Figure 5: Structural causal model for generating data on a target variable $Y$, $p=5$ predictors and the causal graph (a). The structural equations in (b) show how the target $Y$ has a Bernoulli distribution conditional on the causal parents $X_{2}$ and $X_{3}$.
Figure 6: Results across 100 simulations with varying sample sizes from the Model in Figure 5 with a Bernoulli distribution for $Y$ conditional on its causal parents. (a) The percentage of times that the correct model is selected by the proposed causal GLM approach, both under an exhaustive and a stepwise procedure, increases with $n$. (b) The stepwise procedure is computationally more efficient than the exhaustive search.
We repeat the simulation 100 times with varying sample sizes and fit logistic regression models for $Y$ for all possible subsets of the predictors. For each model, a bootstrap test at a $5\%$ significance level is used to evaluate evidence against the null hypothesis of a perfectly dispersed Pearson risk. The results in Figure 6 show how, as with the previous simulation, the accuracy in detecting the true causal model increases with the sample size $n$. Moreover, particularly in the case when p-values are calculated via bootstrap, there is a significant gain in computational efficiency when using the stepwise procedure in place of the exhaustive search.
4 Identifying causal drivers in empirical studies